# Cuntz–Pimsner algebras of partial automorphisms twisted by vector bundles I: Fixed point algebra, simplicity and the tracial state space

### v2

@@ Abstract.
We associate a $C^{*}$-algebra to a partial action of the integers acting on the base space of a vector bundle, using the framework of Cuntz–Pimsner algebras. We investigate the structure of the fixed point algebra under the canonical gauge action, and show that it arises from a continuous field of $C^{*}$-algebras over the base space, generalising results of Vasselli. We also analyse the ideal structure, and show that for a free action, ideals correspond to open invariant subspaces of the base space. This shows that if the action is free and minimal, then the Cuntz–Pimsner algebra is simple. Finally we establish a bijective corrrespondence between tracial states and invariant measures on the base space, thereby calculating part of the Elliott invariant. This generalizes results about the $C^{*}$-algebras associated to homeomorphisms twisted by vector bundles of Adamo, Archey, Forough, Georgescu, Jeong, Strung and Viola.
@@
Funded by GAČR project GF22-07833K and RVO: 67985840. Part of this work was carried out while funded by GAČR project 20-17488Y. This work has been supported by Charles University Research Centre program No. UNCE/24/SCI/022, and by the Charles University project SVV-2023-260721.
@@ 1. Introduction
Of particular interest is the case when the coefficient $C^{*}$-algebra of the $C^{*}$-correspondence is commutative. In addition to generalizing crossed products by minimal homeomorphisms, these Cuntz–Pimsner algebras include the well-studied graph $C^{*}$-algebras, and even the topological graph $C^{*}$-algebras from [28]. They also include crossed products by partial actions of the integers on commutative $C^{*}$-algebras, in the sense of [17].
The study of partial actions in the context of $C^{*}$-algebras was initiated by Exel in [16], where he considered partial actions by the integers. Any crossed product by $\mathbb{Z}$ admits an action of $\mathbb{T}$ (the dual action), however the converse need not hold, that is, a $C^{*}$-algebra with an action of $\mathbb{T}$ need not be isomorphic to a crossed product, even if the action is nice enough. However, in [16], Exel showed that if the action regular [16, Definition 4.4], then it is indeed a crossed product by a partial action. Partial actions have since been extended to more general discrete groups and semigroups and have become an important area of study in their own right [17]. In particular, they have turned out to be an important source of examples in Elliott’s classification program [13, 14].
In [1], Abadie, Eilers, and Exel also considered the question of the structure of circle actions on $C^{*}$-algebras. They defined a crossed product by a Hilbert bimodule, and showed that if a $C^{*}$-algebra $A$ admits a circle action which is semi-saturated [16, Definition 4.1], the $A$ is isomorphic to such a crossed product. They give an example of an algebra admitting such a circle action, whose fixed point subalgebra is a unital commutative $C^{*}$-algebra, but which can be neither a crossed product by a homeomorphism nor a partial automorphism [1, Example 3.3.].
@@
Crossed products by partial automorphisms on compact metric spaces with finite covering dimension also have finite nuclear dimension [20], meaning they are covered by the classification theorem that says simple, separable, nuclear, unital, infinite-dimensional $C^{*}$-algebras which satisfy the UCT and have finite nuclear dimension are classified up to isomorphism by the Elliott invariant (see for example [49]).
We also consider the tracial state space of our algebras. For crossed products by partial automorphisms, there is a bijective correspondence between tracial states of the crossed product by a homeomorphism and invariant measures on the space. The tracial state space is a crucial piece of the Elliott invariant. Indeed, in [15] Elliott was able to extend classification outside the case of real rank to simple inductive limits of interval algebras using $K$-theory along with the tracial state space. (In the case of real rank zero, tracial information is contained in the ordered $K_{0}$-group.) Thanks to classification, we see that isomorphism in certain classes of $C^{*}$-algebras—for example the crossed products of $S^{d}$, $d\geq 3$ and odd integer, by minimal diffeomorphisms, or the simple inductive limits of dimension drop algebras with vanishing $K_{1}$-group—depend only on their tracial state space [43, 25].
@@
In this paper we generalize some of the results of [2] to partial $\mathbb{Z}$-actions on locally compact spaces. The $C^{*}$-correspondences we consider are those with a left action of the form (1). However, we replace the homeomorphism $\alpha$ with a homeomorphism $\theta$ between two open subsets $U$ and $V$ of $X$. Such a partial homeomorphism on $X$ gives rise to a partial $\mathbb{Z}$-action in the sense of [17]. Following [20], usually call $\theta$ a partial automorphism.
We have noted above that for partial crossed products, the coefficient algebra $C(X)$ can be recovered as the fixed point algebra of the canonical gauge action, and that it is a Cartan subalgebra if the action is topologically free. For partial automorphisms twisted by vector bundles, this is only true if the vector bundle is a line bundle. If the vector bundle is higher rank, then one should expect the fixed point algebra to be noncommutative and closely related to the UHF algebra $\mathcal{U}_{n^{\infty}}$, where $n$ is the rank of the vector bundle. This is the case since the fixed point algebra of the Cuntz algebra $\mathcal{O}_{n}$ under the canonical gauge action is given by $\mathcal{U}_{n^{\infty}}$ [9], and $\mathcal{O}_{n}$ can be obtained in our formalism from considering the trivial action and vector bundle on a one-point space. We investigate the structure of the fixed point algebra in our setting in Section 3.
As expected, the ideal structure of the Cuntz–Pimsner algebra is determined by dynamical properties of the action in the same way that this is the case for partial crossed products by the integers. For the tracial state space we obtain the same dichotomy as in [2]: If the vector bundle is a line bundle, meaning that all its fibers are one-dimensional, then the usual bijective correspondence between traces and invariant measures holds. In the other case the Cuntz–Pimsner algebra does not have any traces. In the context of the Elliott classification program this results fits into the well-known dichotomy between stably finite and purely infinite $C^{*}$-algebras.
Further connections to classification will be explored in a second paper: Under suitable conditions on the compact Hausdorff space $X$, it follows from the general theory of Cuntz–Pimsner algebras [29] that the Cuntz–Pimsner algebra of a partial automorphism twisted by a vector bundle is nuclear, separable and satisfies the UCT. In this second paper we will investigate under which assumptions our Cuntz–Pimsner algebras have finite nuclear dimension (and are thus classifiable), and again find that it is analogous to the case of partial crossed products. In future work, we will determine the Elliott-invariant of our algebras. Additionally to the traces, which have been treated in this paper, this includes $K$-theory as well as the pairing between $K$-theory and traces.
@@ 1.1. Summary of the paper
The main results of the paper are contained in Sections 3 and 4. We start in Section 3 by discussing Hilbert $C_{0}(X)$-modules. Since we want to consider locally compact spaces, we need a generalization of the Serre–Swan theorem from the compact case. For this reason we introduce properties of Hilbert $C_{0}(X)$-modules which we call $\sigma$-finitely generated and $\sigma$-projective. By definition, a Hilbert $C_{0}(X)$-module has these properties whenever any restriction to a compact subset of $X$ is finitely generated and projective, respectively. We outline how to show that such modules are exactly the Hilbert $C_{0}(X)$-modules $\Gamma_{0}(\mathcal{V})$ of continuous sections vanishing of infinity of a vector bundle $\mathcal{V}$ over $X$. This is the desired Serre–Swan type theorem.
@@
Section 4 investigates the ideal structure as well as the tracial state space of our Cuntz–Pimsner algebras. Regarding the ideal structure, we show that the known results about (partial) crossed products still hold in our setting. Under the additional assumption of freeness, ideals in the Cuntz–Pimsner algebra correspond to open invariant subsets of the base space $X$, see Theorem 4.16. Thus the Cuntz–Pimsner algebra associated to a free and minimal partial automorphism is simple, regardless of what the vector bundle looks like. Regarding the tracial state space we show that if the vector bundle is not a line bundle then there do not exist any faithful traces, and if the vector bundle is higher rank everywhere then there are no traces at all. In case the vector bundle is rank one, the tracial state space is affinely homeomorphic to the compact convex set of invariant measures on $X$, see Proposition 4.23. This aligns well with results from [2].
@@ 2.1. Hilbert $C^{*}$-modules
Let $A$ be a $C^{*}$-algebra and let $\mathcal{E}$ be a right Hilbert $A$-module. Take a subset $S$ of $\mathcal{E}$. By $\operatorname{span}_{A}$ we denote the $A$-linear span of $S$, namely the submodule of all finite linear combinations of elements of $S$ with coefficients in $A$. We say that $S$ generates $\mathcal{E}$ if $\operatorname{span}_{A}S$ is dense in $\mathcal{E}$. If there exists a finite or countable set generating $\mathcal{E}$ then we call $\mathcal{E}$ finitely generated or countably generated, respectively. If there exists a set $S$ such that $\operatorname{span}_{A}S$ is equal to $\mathcal{E}$, then $\mathcal{E}$ is called algebraically finitely generated. A Hilbert $A$-module is called projective if it is projective as an $A$-module, namely if it is an (algebraic) direct summand in some free $A$-module.
@@
Rank-one operators are dense in $\mathcal{K}(\mathcal{E})$. Hence it is sufficient to show that for all $\xi,\eta\in\mathcal{E}$ the expression $\sum_{i\in F}\theta_{\zeta_{i},\zeta_{i}}\theta_{\xi,\eta}$ can be made arbitrarily close to $\theta_{\xi,\eta}$ by the right choice of a finite subset $F\subset I$. This follows from the inequality
@@ 2.2. Hilbert $C(X)$-modules
An introduction to vector bundles can be found in [24]. We denote a vector bundle over $X$ by $\mathcal{V}=[E,p,X]$ where $E$ is the total space, $X$ is the base space, and $p$ is a surjective continuous map from $E$ to $X$ called the bundle projection. All vector bundles in this work will be complex. As is standard, we require our vector bundles to be locally trivial, meaning that for every point $x\in X$ there exist a neighborhood $U$ of $x$ and a bundle morphism $h$ from $U\times\mathbb{C}^{n_{x}}$ to $p^{-1}(U)$ inducing an isomorphism on each fiber, for some $n_{x}\in\mathbb{N}$. The pair $(U,h)$ is called a chart, and a collection of charts such that the open sets cover $X$ is called an atlas. Note that by definition of a chart, the rank of $\mathcal{V}$ is constant over $U$. More precisely, we have $\operatorname{rank}\mathcal{V}_{y}=n_{x}$ for all $y\in U$.
@@
where (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ) is the standard inner product on $\mathbb{C}^{N_{i}}$.
@@
According to [24, Chapter 5, Theorem 7.4] there exist metric-preserving atlases $\{(h_{i},U_{i})\}$ and $\{(\tilde{h}_{i},\tilde{U}_{i})\}$ associated to $\beta$ and $\tilde{\beta}$, respectively. Let $g_{ij}$ and $\tilde{g}_{ij}$ be systems of transition functions for the atlases. We can assume that the open covers $\{U_{i}\}$ and $\{\tilde{U}_{i}\}$ are the same, otherwise we use the open cover of intersections $U_{i}\cap\tilde{U_{j}}$.
Since both atlases describe the same vector bundle, they are equivalent (see [24, Chapter 5, Theorem 2.7]). Hence there exist functions $r_{i}:U_{i}\to GL(N_{i})$ satisfying $\tilde{g}_{ij}=r_{i}^{-1}g_{ij}r_{j}$. Using this equality, a simple calculation shows $r_{i}r_{i}^{*}g_{ij}=g_{ij}r_{j}r_{j}^{*}$. This implies that we have $P(r_{i}r_{i}^{*})g_{ij}=g_{ij}P(r_{j}r_{j}^{*})$ for all polynomials $P$. Using functional calculus for the positive matrix $r_{i}r_{i}^{*}$ we obtain $|r_{i}|^{-1}g_{ij}=g_{ij}|r_{j}|^{-1}$.
@@ 2.3. $C^{*}$-correspondences and Hilbert bimodules
$\overline{\operatorname{span}}(\varphi_{\mathcal{E}}(A)\mathcal{E})$ is equal to $\mathcal{E}$. Here $\overline{\operatorname{span}}(S)$ denotes the closure of the $A$-linear span of a set $S\subset\mathcal{E}$. The basics about $C^{*}$-correspondences can be found in [29].
@@
Let $A$ be a $C^{*}$-algebra and $\phi:A\to A$ a $\ast$-homomorphism. Equip $A$ with algebra multiplication as right multiplication and the inner product $\langle a,b\rangle_{A}=a^{*}b$. Then $A$ is a right Hilbert $A$-module and $\mathcal{L}(A)$ is isomorphic to $A$. The $\ast$-homomorphism $\phi$ induces a structure map $\varphi_{A}:A\to\mathcal{L}(A)\cong A$ given by $\varphi_{A}(a)b=\phi(a)b$. This yields a $C^{*}$-correspondence over $A$. If $\phi=\operatorname{id}_{A}$ then we denote the resulting $C^{*}$-correspondence by $\mathbf{B}$.
Let $\mathcal{E}$ and $\mathcal{F}$ be $C^{*}$-correspondences over $A$. We denote the internal tensor product of $\mathcal{E}$ and $\mathcal{F}$ by $\mathcal{F}\otimes_{A}\mathcal{E}$. It is a $C^{*}$-correspondence over $A$ with inner product
$\langle\xi_{1}\otimes\eta_{1},\xi_{2}\otimes\eta_{2}\rangle_{\mathcal{E}\otimes\mathcal{F}}\coloneqq\langle\eta_{1},\varphi_{\mathcal{F}}(\langle\xi_{1},\xi_{2}\rangle_{\mathcal{E}})\eta\rangle_{\mathcal{F}}.$
@@
Every Hilbert $A$-bimodule is a $C^{*}$-correspondence over $A$. To see this, define the structure map $\varphi_{\mathcal{E}}$ to act by left multiplication, $\varphi_{\mathcal{E}}(a)\xi=a\xi$. It is shown in [8] that $\varphi_{\mathcal{E}}$ indeed maps into the adjointable operators on $\mathcal{E}$. If we restrict $\varphi_{\mathcal{E}}$ to Katsura’s ideal $J_{\mathcal{E}}$ then it becomes a $\ast$-isomorphism. We can also go the opposite direction: Any $C^{*}$-correspondence such that the structure map restricts to an isomorphism from Katsura’s ideal to the compacts can be turned into a Hilbert $A$-bimodule. Left multiplication is given by $a\xi\coloneqq\varphi_{\mathcal{E}}(a)\xi$, and the inner product by ${}_{\mathcal{E}}\langle\xi,\eta\rangle\coloneqq\varphi_{\mathcal{E}}^{-1}(\theta_{\xi,\eta})$.

### v3

@@ Abstract.
We associate a $C^{*}$-algebra to a partial action of the integers acting on the base space of a vector bundle, using the framework of Cuntz–Pimsner algebras. We investigate the structure of the fixed point algebra under the canonical gauge action, and show that it arises from a continuous field of $C^{*}$-algebras over the base space, generalising results of Vasselli. We also analyse the ideal structure, and show that for a free action, ideals correspond to open invariant subspaces of the base space. This shows that if the action is free and minimal, then the Cuntz–Pimsner algebra is simple. In the case of a line bundle, we establish a bijective corrrespondence between tracial states on the algebra and invariant measures on the base space. This generalizes results about the $C^{*}$-algebras associated to homeomorphisms twisted by vector bundles of Adamo, Archey, Forough, Georgescu, Jeong, Strung and Viola.
@@
Funded by GAČR project GF22-07833K and RVO: 67985840. Part of this work was carried out while funded by GAČR project 20-17488Y. This work has been supported by Charles University Research Centre program No. UNCE/24/SCI/022, and by the Charles University project SVV-2023-260721. The author is currently funded by GAČR project G25-15403K.
@@ 1. Introduction
Of particular interest is the case when the coefficient $C^{*}$-algebra of the $C^{*}$-correspondence is commutative. In addition to generalizing crossed products by homeomorphisms, these Cuntz–Pimsner algebras include the well-studied graph $C^{*}$-algebras, and even the topological graph $C^{*}$-algebras from [28]. They also include crossed products by partial actions of the integers on commutative $C^{*}$-algebras, in the sense of [17].
The study of partial actions in the context of $C^{*}$-algebras was initiated by Exel in [16], where he considered partial actions by the integers. Any crossed product by $\mathbb{Z}$ admits an action of $\mathbb{T}$ (the dual action). The converse need not hold, that is, a $C^{*}$-algebra with an action of $\mathbb{T}$ need not be isomorphic to a crossed product, even if the action is nice enough. However, in [16], Exel showed that if the action is regular [16, Definition 4.4], then it is indeed a crossed product by a partial integer action. This generalised a result by Landstad for global actions [Landstad:1979]. Partial actions have since been extended to more general discrete groups and semigroups and have become an important area of study in their own right [17]. In particular, they have turned out to be an important source of examples in Elliott’s classification program [13, 14].
In [1], Abadie, Eilers, and Exel also considered the question of the structure of circle actions on $C^{*}$-algebras. They defined crossed products by Hilbert bimodules, and showed that a $C^{*}$-algebra is isomorphic to such a crossed product if and only if it admits a circle action which is semi-saturated [16, Definition 4.1]. They give an example of an algebra admitting such a circle action, whose fixed point subalgebra is a unital commutative $C^{*}$-algebra, but which is not a crossed product by a partial automorphism [1, Example 3.3.].
@@
Crossed products by minimal partial automorphisms on compact metric spaces with finite covering dimension also have finite nuclear dimension [20], meaning they are covered by the classification theorem that says simple, separable, nuclear, unital, infinite-dimensional $C^{*}$-algebras which satisfy the UCT and have finite nuclear dimension are classified up to isomorphism by the Elliott invariant (see for example [49]).
A important part of the Elliott invariant is the tracial state space. For crossed products by free partial automorphisms, there is a bijective correspondence between tracial states of the crossed product by a homeomorphism and invariant measures on the space. The tracial state space is a crucial piece of the Elliott invariant. Indeed, in [15] Elliott was able to extend classification outside the case of real rank zero to simple inductive limits of interval algebras using $K$-theory along with the tracial state space. (In the case of real rank zero, tracial information is contained in the ordered $K_{0}$-group.) Thanks to classification, we see that isomorphism in certain classes of $C^{*}$-algebras—for example the crossed products of $S^{d}$, $d\geq 3$ and odd integer, by minimal diffeomorphisms, or the simple inductive limits of dimension drop algebras with vanishing $K_{1}$-group—depend only on their tracial state space [43, 25].
@@
In this paper we generalize some of the results of [2] to partial $\mathbb{Z}$-actions on locally compact spaces. The $C^{*}$-correspondences we consider are those with a left action of the form (1). However, we replace the homeomorphism $\alpha$ with a homeomorphism $\theta$ between two open subsets $U$ and $V$ of $X$. Such a partial homeomorphism on $X$ gives rise to a partial $\mathbb{Z}$-action in the sense of [17]. We usually call $\theta$ a partial automorphism.
We have noted above that for partial crossed products, the coefficient algebra $C(X)$ can be recovered as the fixed point algebra of the canonical gauge action, and that it is a Cartan subalgebra if the action is topologically free. For partial automorphisms twisted by vector bundles, this is only true if the vector bundle is a line bundle. If the vector bundle is higher rank, then one should expect the fixed point algebra to be noncommutative and closely related to the UHF algebra $M_{n^{\infty}}$, where $n$ is the rank of the vector bundle. This is the case since the fixed point algebra of the Cuntz algebra $\mathcal{O}_{n}$ under the canonical gauge action is given by $M_{n^{\infty}}$ [9], and $\mathcal{O}_{n}$ can be obtained in our formalism from considering the trivial action and vector bundle on a one-point space. We investigate the structure of the fixed point algebra in our setting in Section 3.
As expected, for a free action the ideal structure of the Cuntz–Pimsner algebra is determined by dynamical properties of the action in the same way that this is the case for partial crossed products by the integers. For the tracial state space we generalise results from [2], again in the case of a free action: If the vector bundle is a line bundle, meaning that all its fibers are one-dimensional, then the usual bijective correspondence between traces and invariant measures holds. In particular, those Cuntz–Pimsner algebras are stably finite.
Further connections to classification will be explored in a second paper: Under suitable conditions on the compact Hausdorff space $X$, it follows from the general theory of Cuntz–Pimsner algebras [29] that the Cuntz–Pimsner algebra of a partial automorphism twisted by a vector bundle is nuclear, separable and satisfies the UCT. In this second paper we will investigate under which assumptions our Cuntz–Pimsner algebras have finite nuclear dimension (and are thus classifiable), and again find that it is analogous to the case of partial crossed products. In future work, we will determine the Elliott-invariant of our algebras. Additionally to the traces, this includes $K$-theory as well as the pairing between $K$-theory and traces.
@@ 1.1. Summary of the paper
The main results of the paper are contained in Sections 3 and 4. We start in Section 3 by discussing Hilbert $C_{0}(X)$-modules. Since we want to consider locally compact spaces, we need a generalization of the Serre–Swan theorem from the compact case. For this reason we introduce properties of Hilbert $C_{0}(X)$-modules which we call $\sigma$-finitely generated and $\sigma$-projective. By definition, a Hilbert $C_{0}(X)$-module has these properties whenever any restriction to a compact subset of $X$ is finitely generated and projective, respectively. We outline how to show that such modules are exactly the Hilbert $C_{0}(X)$-modules $\Gamma_{0}(\mathcal{V})$ of continuous sections vanishing at infinity of a vector bundle $\mathcal{V}$ over $X$. This is the desired Serre–Swan type theorem.
@@
Section 4 investigates the ideal structure as well as the tracial state space of our Cuntz–Pimsner algebras. Regarding the ideal structure, we show that the known results about (partial) crossed products still hold in our setting. Under the additional assumption of freeness, ideals in the Cuntz–Pimsner algebra correspond to open invariant subsets of the base space $X$, see Theorem 4.16. Thus the Cuntz–Pimsner algebra associated to a free and minimal partial automorphism is simple, regardless of what the vector bundle looks like. We show that if the vector bundle is rank one, then the tracial state space is affinely homeomorphic to the compact convex set of invariant measures on $X$, see Proposition 4.23. This aligns well with results from [2].
@@ 2.1. Hilbert $C^{*}$-modules
Let $A$ be a $C^{*}$-algebra and let $\mathcal{E}$ be a right Hilbert $A$-module. Take a subset $S$ of $\mathcal{E}$ and write
$\operatorname{span}_{A}S=\{\sum_{i=1}^{n}s_{i}a_{i}|n\in\mathbb{N},s_{1},...,s_{n}\in S,a_{1},...,a_{n}\in A\}$
for the $A$-linear span of $S$. We say that $S$ generates $\mathcal{E}$ if $\operatorname{span}_{A}S$ is dense in $\mathcal{E}$. If there exists a finite or countable set generating $\mathcal{E}$ then we call $\mathcal{E}$ finitely generated or countably generated, respectively. If there exists a set $S$ such that $\operatorname{span}_{A}S$ is equal to $\mathcal{E}$, then $\mathcal{E}$ is called algebraically finitely generated. A Hilbert $A$-module is called projective if it is projective as an $A$-module, namely if it is an (algebraic) direct summand in some free $A$-module.
@@
Rank-one operators densely span $\mathcal{K}(\mathcal{E})$. Hence it is sufficient to show that for all $\xi,\eta\in\mathcal{E}$ the expression $\sum_{i\in F}\theta_{\zeta_{i},\zeta_{i}}\theta_{\xi,\eta}$ can be made arbitrarily close to $\theta_{\xi,\eta}$ by the right choice of a finite subset $F\subset I$. This follows from the inequality
@@ 2.2. Hilbert $C(X)$-modules
An introduction to vector bundles can be found in [24]. We denote a vector bundle over $X$ by $\mathcal{V}=[E,p,X]$ where $E$ is the total space, $X$ is the base space, and $p$ is a surjective continuous map from $E$ to $X$ called the bundle projection. All vector bundles in this work will be complex.
@@
As is standard, we require our vector bundles to be locally trivial, meaning that for every point $x\in X$ there exist a neighborhood $U$ of $x$ and a bundle morphism $h$ from $U\times\mathbb{C}^{n_{x}}$ to $p^{-1}(U)$ inducing an isomorphism on each fiber, for some $n_{x}\in\mathbb{N}$. The pair $(U,h)$ is called a chart, and a collection of charts such that the open sets cover $X$ is called an atlas. Note that by definition of a chart, the rank of $\mathcal{V}$ is constant over $U$. More precisely, we have $\operatorname{rank}\mathcal{V}_{y}=n_{x}$ for all $y\in U$.
@@
where (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ) is the standard inner product on $\mathbb{C}^{N_{i}}$. Metric preserving atlases have unitary transition functions.
@@
According to [24, Chapter 5, Theorem 7.4] there exist metric-preserving atlases $\{(h_{i},U_{i})\}$ and $\{(\tilde{h}_{i},\tilde{U}_{i})\}$ associated to $\beta$ and $\tilde{\beta}$, respectively. Let $g_{ij}$ and $\tilde{g}_{ij}$ be systems of unitary transition functions for the atlases. We can assume that the open covers $\{U_{i}\}$ and $\{\tilde{U}_{i}\}$ are the same, otherwise we use the open cover of intersections $U_{i}\cap\tilde{U_{j}}$.
Since both atlases describe the same vector bundle, they are equivalent (see [24, Chapter 5, Theorem 2.7]). Hence there exist functions $r_{i}:U_{i}\to GL(N_{i})$ satisfying $\tilde{g}_{ij}=r_{i}^{-1}g_{ij}r_{j}$. Using this equality and the fact that $g_{ij}^{*}=g_{ij}^{-1}=g_{ji}$, we obtain
$r_{i}r_{i}^{*}g_{ij}=r_{i}(g_{ij}^{*}r_{i})^{*}=r_{i}(g_{ji}r_{i})^{*}=r_{i}(r_{j}\tilde{g}_{ji})^{*}=r_{i}\tilde{g}_{ij}r_{j}^{*}=g_{ij}r_{j}r_{j}^{*}.$
This implies $P(r_{i}r_{i}^{*})g_{ij}=g_{ij}P(r_{j}r_{j}^{*})$ for all polynomials $P$. Using functional calculus for the positive matrix $r_{i}r_{i}^{*}$ we obtain $|r_{i}|^{-1}g_{ij}=g_{ij}|r_{j}|^{-1}$.
@@ 2.3. $C^{*}$-correspondences and Hilbert bimodules
$\overline{\operatorname{span}_{A}}(\varphi_{\mathcal{E}}(A)\mathcal{E})$ is equal to $\mathcal{E}$. Here $\overline{\operatorname{span}_{A}}(S)$ denotes the closure of the $A$-linear span of a set $S\subset\mathcal{E}$. The basics about $C^{*}$-correspondences can be found in [29].
@@
Let $A$ be a unital $C^{*}$-algebra and $\phi:A\to A$ a $\ast$-homomorphism. Equip $A$ with algebra multiplication as right multiplication and the inner product $\langle a,b\rangle_{A}=a^{*}b$. Then $A$ is a right Hilbert $A$-module and $\mathcal{K}(A)$ is isomorphic to $A$. The $\ast$-homomorphism $\phi$ induces a structure map $\varphi_{A}:A\to\mathcal{K}(A)$ given by $\varphi_{A}(a)b=\phi(a)b$. This yields a $C^{*}$-correspondence over $A$. If $\phi=\operatorname{id}_{A}$ then we denote the resulting $C^{*}$-correspondence by $\mathbf{A}$.
Let $\mathcal{E}$ and $\mathcal{F}$ be $C^{*}$-correspondences over $A$. We denote the internal tensor product of $\mathcal{E}$ and $\mathcal{F}$ by $\mathcal{E}\otimes_{A}\mathcal{F}$. It is a $C^{*}$-correspondence over $A$ with inner product
$\langle\xi_{1}\otimes\eta_{1},\xi_{2}\otimes\eta_{2}\rangle_{\mathcal{E}\otimes\mathcal{F}}\coloneqq\langle\eta_{1},\varphi_{\mathcal{F}}(\langle\xi_{1},\xi_{2}\rangle_{\mathcal{E}})\eta_{2}\rangle_{\mathcal{F}}.$
@@
Every Hilbert $A$-bimodule is a $C^{*}$-correspondence over $A$. To see this, define the structure map $\varphi_{\mathcal{E}}$ to act by left multiplication, $\varphi_{\mathcal{E}}(a)\xi=a\xi$. It is shown in [8] that $\varphi_{\mathcal{E}}$ indeed maps into the adjointable operators on $\mathcal{E}$. If we restrict $\varphi_{\mathcal{E}}$ to Katsura’s ideal $J_{\mathcal{E}}$ then it becomes a $\ast$-isomorphism onto $\mathcal{K}(\mathcal{E})$. We can also go the opposite direction: Any $C^{*}$-correspondence such that the structure map restricts to an isomorphism from Katsura’s ideal to the compacts can be turned into a Hilbert $A$-bimodule. Left multiplication is given by $a\xi\coloneqq\varphi_{\mathcal{E}}(a)\xi$, and the inner product by ${}_{\mathcal{E}}\langle\xi,\eta\rangle\coloneqq\varphi_{\mathcal{E}}^{-1}(\theta_{\xi,\eta})$.